OpenStudy (ndsncjas):

Write the equation of the line in Standard Form that passes through the points (3, 2) and (5, 8). Question 8 options: y = 7x + 3 -3x + y = 7 3x + y = -7 -3x + y = -7

OpenStudy (jiteshmeghwal9):

(y-y1)=m(x-x1) * @Kevin

OpenStudy (kevin):

lol.. yeah that's what I mean xD

OpenStudy (kevin):

m = (y2-y1)/(x2-x1) y - y1 = m(x - x1)